OpenStudy (anonymous):

Find the equation of the tangent line to f(x)=x^2+4 at x=5

OpenStudy (tkhunny):

Have you considered a 1st Derivative?

OpenStudy (anonymous):

ok I got 10 to

OpenStudy (tkhunny):

"10" is NOT your answer. Please read the problem statement and please don't do other peoples' homework for them.

OpenStudy (anonymous):

Where did I go wrong derivatives are a little confusing to me so cn you help me throught this.

OpenStudy (tkhunny):

You have the slope of the tangent line at x = 5. Now, you need the equation of the tangent line. What information are you missing and how can you find it?

OpenStudy (anonymous):

Don;t I have to find the slope

OpenStudy (anonymous):

oh no sorry I have the slope so then don't I have to find the b in the slope equation

OpenStudy (anonymous):

what would my y be in the y=mx+b equation

OpenStudy (tkhunny):

@farmergirl411 Very good. That is the slope-intercept form. Sadly, we have no way of finding the y-intercept. Do you know the point-slope form? @mathematician1 The problem statement calls for the EQUATION. You stopped at the SLOPE. This is NOT a correct response.

OpenStudy (anonymous):

How do I do the point slope form?

OpenStudy (tkhunny):

We need a point. The problem statememt gives us one. We have x = 5, the we get f(5) = 5^2 + 4 = 29 and we have a point, (5,29). Do you see that?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Perfect. All that is left is to substitute into the point-slope form of a line. (y-29) = 10(x-5) That is all there is to it. Of course, someone may wish that you first out it into Slope-Intercept form. Just solve for y and that will be done.

OpenStudy (anonymous):

so the would the equation be y=10-21

OpenStudy (anonymous):

sorry y=10x-21?

OpenStudy (tkhunny):

Awesome! Any part of that seem confusing, still?

OpenStudy (anonymous):

no but I am just curios where did the 10 come from in the part that you posted and then you said to solve for x.

OpenStudy (tkhunny):

f'(x) evaluated at x = 5 This gives the slope of the tangent line at the given point. This also gives the slope of the line in the point-slope form. The two 10s are one in the same.

OpenStudy (anonymous):

ok thanks for the explanation and your help.